OpenStudy (anonymous):

What is the surface area of a rectangular prism with a base length of 7 yd, a base width of 5 yd, and a height of 6 yd?

OpenStudy (anonymous):

A=2(wl+hl+hw)=214

OpenStudy (anonymous):

A=2(wl+hl+hw)=2·(5·7+6·7+6·5)=214